OpenStudy (flvskidd):

Amanda has a job transporting soft drinks by truck. Her truck is filled with cans that weigh 14 ounces each and bottles that weigh 70 ounces each. There is a combined total of 860 cans and bottles in her truck. Let x be the number of 14 -ounce cans in her truck. Write an expression for the combined total weight (in ounces) of the cans and bottles in her truck.

MARC (marc_d):

Hmm,what do u think? @FLVSKidd

MARC (marc_d):

\[Let~x~be~the~number~of~14~ounce~cans~in~her~truck\]\[Let~y~be~the~number~of~70~ounce~bottles~in~her~truck\]\[Combined~total~of~860~cans~and~bottles~in~her~truck.\] Using the given information..u can start to form an expression

MARC (marc_d):

\[14x+70y=860\]

OpenStudy (mathstudent55):

@MARC_D wrote: "Using the given information..u can start to form an expression 14x+70y=860" The above expression is incorrect.

OpenStudy (mathstudent55):

There is a combined total of 860 cans and bottles. number of cans + number of bottles = 860 The problem instructs us to use x to represent the number of cans, so we can write x + number of bottles = 860 Can you solve the above equation for the number of bottles?

OpenStudy (mathstudent55):

Subtract x from both sides to get number of bottles = 860 - x Now we know this: number of cans = x number of bottles = 860 - x

OpenStudy (mathstudent55):

Each can weights 14 oz, and each bottle weighs 70 oz. The number of cans times 14 oz is the total weight of the cans. The number of bottles times 70 oz is the total weight of the bottles. The sum of the two amounts above is the total combined weight of cans and bottles.